the maxium grade of the railway at pikes peak is 25 percent. find the angle of inclination of the railway at this point
Answer by stanbon(75887)

the maxium grade of the railway at pikes peak is 25 percent. find the angle of inclination of the railway at this point
-----
25% grade means the road rises 25 ft. for every 100 ft forward.
------
tan(theta) = 25/100 = 1/4
-------
arctan(1/4) = 14.04 degrees
==================================
Cheers,
Stan H.